OpenStudy (anonymous):

Solve the following equation by clearing fractions. Please show all of your work. x/6+2=4/5-x

jhonyy9 (jhonyy9):

(x+12)/6 =(4-5x)/5 5(x-12)=6(4-5x) 5x-60=24-30x 35x=84 x=84/35